# Distribution System Reconfiguration to Mitigate Load Altering Attacks via Stackelberg Games

### v5

@@ II-C1 Branch Flow Model
where $v_{\pi_{i}}-v_{i}=z_{\pi_{i},i}\ I_{\pi_{i},i},$ $s_{\pi_{i},i}=v_{\pi_{i}}\ I_{\pi_{i},i}^{*}$. Note that superscript $(\cdot)^{*}$ denotes the conjugate of a complex number and $\pi_{i}$ represents the parent node of the bus $i$. Furthermore, terms $s_{i,k}$ and $I_{\pi_{i},i}$ illustrate the apparent power and current running between two nodes, $z_{\pi_{i},i}$ is the impedance of the line, and $s_{i}$ is the apparent power in bus $i$.
@@ II-C2 Linearized Distribution Flow
Linearized distribution flow (LinDistFlow) simplifies the branch flow model described in (3) by neglecting branch power losses [33] and is widely adopted in several DN studies. The power flow equations under this model are given by $\sum_{k:(i,k)\in\mathcal{L}}p_{ik}=p_{ji}-p_{i}$, $\sum_{k:(i,k)\in\mathcal{L}}q_{ik}=q_{ji}-q_{i}$, as active and reactive power balances and $v^{2}_{i}=v_{j}^{2}-2(r_{ij}p_{ji}+x_{ij}q_{ji})$ the equation for finding subsequent voltage profile. Note that $p_{ik}$ and $q_{ik}$ are active and reactive power running the corresponding line, $p_{i}$ and $q_{i}$ are active and reactive loads in bus $i$, while $r_{ij}$ and $x_{ij}$ are resistance and reactance of lines.
@@ II-D Reconfiguration of DN
Network reconfiguration involves modifying the DN’s topology by adjusting the open/closed state of its switches [22]. In this subsection, a set of MILP optimization constraints is introduced primarily to determine the configuration that maintains the nodal voltages closest to their nominal values. We modify the formulation in [34] to accommodate the ZP approximation of the loads to capture its voltage dependency. We also change the power flow model to the LinDistFlow. As we show in Section V, these approximations [31, 33] result in a MILP the network reconfiguration problem and provide significant computational speed-ups compared to the MISOCP.
@@
$\sum_{j\in\mathcal{N}}p_{ji}-p_{ij}=p_{i}^{ZP},\sum_{j\in\mathcal{N}}q_{ji}-q_{ij}=q_{i}^{ZP},\hskip 8.53581pti\in\mathcal{N}\setminus\mathcal{N}^{f}$ (6)
$s^{zp}_{i}(v_{i})=p^{l}_{i}(\alpha^{\prime}_{p}+\gamma^{\prime}_{p}v_{i}^{2})+jq^{l}_{i}(\alpha^{\prime}_{q}+\gamma^{\prime}_{q}v^{2}_{i}),$ (7)
@@
Here $b_{ij}$ represents the elements of the connectivity matrix of the grid, $\mathcal{N}^{f}$ is the set of substation bus(es), $M$ is the disjunctive parameter, $p_{i}^{f}$ and $q_{i}^{f}$ are active and reactive power flows from substation bus $i$, $\underline{v_{i}}$ and $\overline{v_{i}}$ are voltage limits, and $\hat{v}$ is an auxiliary variable. Equations (4) - (6) represent the power flow constraints, (7) is ZIP load constraint, and (8) - (11) are voltage constraints. The auxiliary variable of $\hat{v}$ makes the optimization follow disciplined convex programming rules in Python. Additionally, we consider all the lines to have switches.
II-E Load-Altering Attack Model
Some major manufacturers of high-wattage IoT-controllable devices have acknowledged the presence of security vulnerabilities in their products [11]. The concept presented in the LAAs involves attackers leveraging these vulnerabilities to transform a group of such devices into bots and toggle them on and off. This coordinated action is intended to disrupt the stability of the system. Based on this, to power balance equations change into
@@ II-E Load-Altering Attack Model
where $x$ is a short-hand notation, with $x=p$ for active power and $x=q$ for reactive power. Herein, $p^{a}$ and $q^{a}$ are additional active and reactive powers change due to the LAA in bus $i$ and $\mathcal{N}^{a}$ is the set of attacked buses. In the following, based on the provided models and context, we first analyze the impact of LAA on DNs. Then, we propose a game-theoretic mitigation scheme for these attacks.
@@ III Effects of LAA on DNs
In this section, we analyze the impact of LAA on DNs. Our objective is to derive closed-form expressions for the voltage profile of the network with voltage-dependant loads under LAA and for the minimum compromised devices required to cause nodal voltage safety violations. Note that grid’s voltage under LAA can also be computed by solving the power flow equation (3) through an iterative approach such as the backward-forward sweep (BFS) technique. However, unlike closed-form equations (which we derive in this section), the iterative techniques does not yield analytical insights into the impact of LAAs. Furthermore, the closed-form expressions obtained in this section are essential in the design of the defense strategies to mitigate LAAs.
@@ III-A1 Without LAA
$v_{k}=\sqrt{v_{1}^{2}-2\sum_{i\in\mathcal{D}_{k}}\left(r_{\pi_{i},i}p^{zp}_{\pi_{i},i}+x_{\pi_{i},i}q^{zp}_{\pi_{i},i}\right)},$ (13)
where $p_{\pi_{i},i}^{zp}=p_{\pi_{i},i}(\alpha^{\prime}_{p}+\gamma^{\prime}_{p}v_{i}^{2}),$ $q_{\pi_{i},i}^{zp}=q_{\pi_{i},i}(\alpha^{\prime}_{q}+\gamma^{\prime}_{q}v_{i}^{2}).$ Next, we perform a variable change $(u_{k}=v_{k}^{2}),$ which results in a set of linear equations, which can be written in matrix form as $\textbf{U}_{(N-1)\times 1}=\Omega_{(N-1)\times N}\begin{bmatrix}1\\ \textbf{U}\end{bmatrix}_{N\times 1}$, where U is the vector of squares of voltages, and $\Omega_{(n-1)\times n}$ is the matrix with entries:
@@
[…]
where $2\leq i,k\leq N$ ($N$ is number of buses). We re-write the system of linear equations as $(\textbf{I}_{(N-1)\times(N-1)}-\Omega^{\prime}_{(N-1)\times(N-1)})\ \textbf{U}_{(N-1)\times 1}=\Omega^{\prime\prime}_{(N-1)\times 1}$, in which $\Omega^{\prime\prime}_{(N-1)\times 1}=[\omega_{2,k}]$, $\Omega^{\prime}_{(N-1)\times(N-1)}=[\omega_{i,k}]$ for $i=\{3,4,...,N\}$, and $k\in\mathcal{N}$.
@@ III-A2 With LAA
Here, we analyze the voltage profile of the network under LAA. For this, we integrate the introduced LAA in Section II-E into (13) and obtain
@@
in which $\Delta_{k}=\sum_{i\in\mathcal{D}_{k}}\left(r_{\pi_{i},i}p^{zp}_{\pi_{i}^{,}i}+x_{\pi_{i},i}q^{zp}_{\pi_{i},i}\right)$,
$r_{k,a}=\sum_{i\in\{\mathcal{D}_{a}\cap\mathcal{D}_{k}\}}r_{\pi_{i},i},$ and $x_{k,a}=\sum_{i\in\{\mathcal{D}_{a}\cap\mathcal{D}_{k}\}}x_{\pi_{i},i}$. This change results in a new set of coefficient matrices. To calculate the attacked system’s square of voltages vector $(\textbf{U}^{A})$, we solve $\textbf{U}^{A}_{(N-1)\times 1}=\Omega^{A}_{(N-1)\times N}\begin{bmatrix}1\\ \textbf{U}^{A}\end{bmatrix}_{N\times 1}$. To obtain the coefficients matrices, (12) is dragged into LinDistFlow as the ZP model is imposed on them. The final results are $\omega_{i,k}^{A}=\omega_{i,k}+\omega_{i,k}^{a}$, for $i\geq 2$ and $k\geq 1$; $\Omega^{A}_{(N-1)\times N}=[\Omega^{A^{\prime\prime}}_{(N-1)\times 1}\ \Omega^{A^{\prime}}_{(N-1)\times(N-1)}]$, in which for $i\geq 2$ and $k\geq 2$ and $\omega_{i,1}^{a}=\sum_{c\in\{\mathcal{D}_{i}\cap\mathcal{D}_{a}\}}-2p_{a}^{A_{0}}\alpha^{\prime}_{p}r_{\pi_{c},c}-2q_{a}^{A_{0}}\alpha^{\prime}_{q}x_{\pi_{c},c}$ , $\omega_{i,k}^{a}=\begin{cases}-2r_{\pi_{i},i}p^{a}\gamma^{\prime}_{p}-2x_{\pi_{i},i}q^{a}\gamma^{\prime}_{q},\ \ \text{if}\hskip 5.69046pti\in\mathcal{D}_{a}\\ \omega_{\pi_{i},k}^{A},\ \ \text{otherwise}.\end{cases}$
@@
Intuitively, higher power flows in distribution lines cause bigger voltage drops. The additional demand caused by LAAs in deeper nodes leads to higher load flows in more lines (to convey the power from the root to the LAA location). As a result, these LAAs cause more drastic voltage drops.
@@ III-B Analytical Insights into the Attack Impact
Next, we use the closed-form expressions obtained above to derive analytical insights into the attack impact. First, note that when there is an attack in a leaf bus (the last bus of each branch), $r_{k,a}$ and $x_{k,a}$ have the highest possible values. As a result, the voltage drop resulted from $(-2p^{a}r_{k,a}-2q^{a}x_{k,a})$ in (16) is higher and obtained voltages shrink. In conclusion, attacks on the leaf buses yield the most detrimental effects.
While this result is somewhat expected for DNs, we further obtain the minimum number of attacked devices, which leads to voltage safety violations. We call such a threat the “critical attack”. For this, we consider the voltage of the leaf bus as a known variable ($v_{th}$). The new unknown variable is $p^{a}$, and based on the attacked device type, we can find $q^{a}$ via $q^{a}=\frac{q_{d}}{p_{d}}\ p^{a}$. So the new set of coefficients for obtaining voltages of buses except for the leaf one and the active power of the critical attack is forming $\Omega^{d}$ ($\Omega^{d}=[\Omega^{d^{\prime\prime}}\Omega^{d^{\prime}}]$) in which
$\omega_{i,1}^{d}$ =$=\begin{cases}\omega_{i,1}+v_{th}^{2}\omega_{i,a},\ \ \text{if}\ i\neq a,\\ \omega_{i,1}+v_{th}^{2}(\omega_{i,a}-1),\ \ \text{if}\ i=a,\end{cases}$ (17)
$\omega_{i,k}^{d}$ =$=\begin{cases}\sum_{c\in\mathcal{D}_{i}}-2r_{\pi_{c},c}\alpha^{\prime}_{p}-2\frac{q_{d}}{p_{d}}x_{\pi_{c},c}\alpha^{\prime}_{q},\ \ \\ \quad\quad\quad\ \text{if}\ k=a,i\in\mathcal{D}_{a},\\ \omega_{\pi_{i},k}^{d},\ \ \text{if}\ k=a,i\notin\mathcal{D}_{a},\\ \omega_{i,k},\ \ \text{otherwise}.\end{cases}$ (18)
Hence, we can solve the linear system of equations given by $(\textbf{I}^{a}-\Omega^{d^{\prime}})\textbf{X}=\Omega^{d^{\prime\prime}}$, in which X is a vector with the same dimension as U where all elements of it are the same as U except for one in which X contains $p^{a}$ instead of $u_{a}$ (since we already know $u_{a}=v_{th}^{2}$). Additionally, $\textbf{I}^{a}$ is the identity matrix except for the element $(a,a)$ which equals 00. This gives the bus voltages as well as the required $p^{a}$. Then, we can find the number of required devices in bus $n$ using $c(n)=\frac{p^{a}}{p_{d}}$.
The closed-form expressions provided in this section are not only used for identifying the worst effect of the LAAs, but also in the following sections, they inspire our mitigation method to find the optimal defensive action. Additionally, the introduced “critical attack” has been implemented in finding the optimal action of the strategic attacker later in this paper.

### v6

@@ II-C1 Branch Flow Model
where $v_{\pi_{i}}-v_{i}=z_{\pi_{i},i}\ I_{\pi_{i},i},$ $s_{\pi_{i},i}=v_{\pi_{i}}\ I_{\pi_{i},i}^{*}$. Note that superscript $(\cdot)^{*}$ denotes the conjugate of a complex number.
@@ II-C2 Linearized Distribution Flow
Linearized distribution flow (LinDistFlow) simplifies the branch flow model described in (3) by neglecting branch power losses [33] and is widely adopted in several DN studies. The power flow equations under this model are given by $\sum_{k:(i,k)\in\mathcal{L}}p_{ik}=p_{ji}-p_{i}$, $\sum_{k:(i,k)\in\mathcal{L}}q_{ik}=q_{ji}-q_{i}$, as active and reactive power balances and $v^{2}_{i}=v_{j}^{2}-2(r_{ij}p_{ji}+x_{ij}q_{ji})$ the equation for finding subsequent voltage profile.
@@ II-D Reconfiguration of DN
Reconfiguration of DNs is a well-studied topic in which the topology of the networks is changed by turning existing line switches on/off [34]. The existence of switches in lines, which can also be used to isolate potential faults in lines, provides a paramount number of possible configurations for the grid. This typically occurs to reduce power loss, balance the network, rectify the voltage profile, enhance network restoration, and improve network reliability [28, 35]. The fact that the operator can perform the changes in topology swiftly, only by sending on/off commands to switches, makes this capability a great fit for responding to a potential LAA.
In this subsection, a set of mixed-integer linear programming (MILP) constraints is introduced primarily to determine the configuration that maintains the nodal voltages closest to their nominal values. We integrate the ZP approximation model into the power flow constraints (which are based on LinDistFlow model) to capture their voltage dependency. As we show in Section V, these approximations result in formulating an MILP for the network reconfiguration problem and provide significant computational speed-ups compared to the mixed-integer second-order cone programming (MISOCP) one.
@@
$\sum_{j\in\mathcal{N}}p_{ji}-p_{ij}=p_{i}^{\text{zp}},\sum_{j\in\mathcal{N}}q_{ji}-q_{ij}=q_{i}^{\text{zp}},\hskip 8.53581pti\in\mathcal{N}\setminus\mathcal{N}^{f}$ (6)
$s^{\text{zp}}_{i}(v_{i})=p^{l}_{i}(\alpha^{\prime}_{p}+\gamma^{\prime}_{p}v_{i}^{2})+jq^{l}_{i}(\alpha^{\prime}_{q}+\gamma^{\prime}_{q}v^{2}_{i}),$ (7)
@@
Equations (4) - (6) represent the power flow constraints, (7) is ZIP load constraint, and (8) - (11) are voltage constraints. The auxiliary variable of $\hat{v}$ makes the optimization follow disciplined convex programming rules in Python. Additionally, we consider all the lines to have switches.
II-E Inverter-Based DER
Inverter-based DERs can adjust the ratio of active to reactive power in their output. We use the following constraint to model such generations in DNs:
$\sqrt{p_{\text{der},i}^{2}+q_{\text{der},i}^{2}}\leq\overline{s_{\text{der},i}}$ (12)
Note that, to integrate DERs into the power flow model, we consider them as negative loads.
II-F Load-Altering Attack Model
Some major manufacturers of high-wattage IoT-controllable devices have acknowledged the presence of security vulnerabilities in their products. The concept presented in the LAAs involves attackers leveraging these vulnerabilities to transform a group of such devices into bots and toggle them on and off. This coordinated action is designed to disrupt the system’s stability. Based on this, power balance equations change into
@@ II-F Load-Altering Attack Model
where $x$ is a short-hand notation, with $x=p$ for active power and $x=q$ for reactive power. LAA differs from normal load deviations in cause, timing, scale, and detectability. Unlike random forecasting errors or unforeseen events, LAA is adversary-driven, strategically coordinated, and deliberately timed to exploit vulnerable grid nodes, often exceeding typical variability bounds. Their targeted and synchronous nature makes them more disruptive and distinguishable from ordinary, statistically distributed load swings.
In the following, based on the provided models and context, we first analyze the impact of LAA on DNs. Then, we propose a game-theoretic mitigation scheme.
II-F1 EVCS Load Altering Attack
As detailed in [36], multiple devices, including EV chargers, air conditioners, heat pumps, inverters, and LED TVs, could be compromised as LAA targets. To provide an example of the attack path, we use the specific case of LAA, in which the attacker targets EVCS. Several common vulnerabilities and exploits (CVEs) have been identified in EV chargers [9]. For instance, CVE-2024-43659 details a vulnerability in a commercial AC EV charger that allows an attacker to obtain default credentials from the firmware. According to this CVE, access to a charger’s firmware exposes the file containing the default credentials shared across all chargers of this model. This is a critical issue because many users do not change the default password, granting an attacker high-privilege access. This vulnerability provides the opportunity to control a large number of these devices by executing false commands. Synchronized execution of these commands on chargers in a concentrated region results in a successful LAA.
@@ III Effects of LAA on DNs
In this section, we analyze the impact of LAA on DNs. Our objective is to derive closed-form expressions for the voltage profile of the network with voltage-dependent loads under LAA and for the minimum compromised devices required to cause nodal voltage safety violations. Note that the grid’s voltage under LAA can also be computed by solving the power flow equation (3) through an iterative approach such as the backwards-forward sweep (BFS) technique. However, unlike closed-form equations (which we derive in this section), the iterative techniques do not yield analytical insights into the impact of LAA. Furthermore, the closed-form expressions obtained in this section are crucial for designing defense strategies to mitigate LAA.
@@ III-A1 Without LAA
$v_{k}=\sqrt{v_{1}^{2}-2\sum_{i\in\mathcal{D}_{k}}\left(r_{\pi_{i},i}p^{\text{zp}}_{\pi_{i},i}+x_{\pi_{i},i}q^{\text{zp}}_{\pi_{i},i}\right)},$ (14)
where $p_{\pi_{i},i}^{\text{zp}}=p_{\pi_{i},i}(\alpha^{\prime}_{p}+\gamma^{\prime}_{p}v_{i}^{2}),$ $q_{\pi_{i},i}^{\text{zp}}=q_{\pi_{i},i}(\alpha^{\prime}_{q}+\gamma^{\prime}_{q}v_{i}^{2}).$ Next, we perform a variable change $(u_{k}=v_{k}^{2}),$ which results in a set of linear equations, which can be written in matrix form as $\textbf{U}_{(N-1)\times 1}=\Omega_{(N-1)\times N}\begin{bmatrix}1\\ \textbf{U}\end{bmatrix}_{N\times 1}$, where U is the vector of squares of voltages, and $\Omega_{(n-1)\times n}$ is the matrix with entries:
@@
$\omega_{i,k}=\begin{cases}\sum_{c=2}^{i}-2r_{\pi_{c},c}p_{k}^{0}\gamma^{\prime}_{p}-2x_{\pi_{c},c}q_{k}^{0}\gamma^{\prime}_{q},\ \ \text{if}\hskip 5.69046pti\in\mathcal{D}_{k}\\ \omega_{\pi_{i},k},\ \ \text{otherwise},\end{cases}$ (16)
where $2\leq i,k\leq N$ (NN is number of buses). We rewrite the system of linear equations as $(\textbf{I}_{(N-1)\times(N-1)}-\Omega^{\prime}_{(N-1)\times(N-1)})\ \textbf{U}_{(N-1)\times 1}=\Omega^{\prime\prime}_{(N-1)\times 1}$, in which $\Omega^{\prime\prime}_{(N-1)\times 1}=[\omega_{2,k}]$, $\Omega^{\prime}_{(N-1)\times(N-1)}=[\omega_{i,k}]$ for $i=\{3,4,...,N\}$, and $k\in\mathcal{N}$.
@@ III-A2 With LAA
Here, we analyze the voltage profile of the network under LAA. For this, we integrate the introduced LAA in Section II-F into (14) and obtain:
@@
in which $\Delta_{k}=\sum_{i\in\mathcal{D}_{k}}\left(r_{\pi_{i},i}p^{\text{zp}}_{\pi_{i}^{,}i}+x_{\pi_{i},i}q^{\text{zp}}_{\pi_{i},i}\right)$, $r_{k,a}=\sum_{i\in\{\mathcal{D}_{a}\cap\mathcal{D}_{k}\}}r_{\pi_{i},i},$ and $x_{k,a}=\sum_{i\in\{\mathcal{D}_{a}\cap\mathcal{D}_{k}\}}x_{\pi_{i},i}$. Further details of obtaining and the rationale behind it are provided in Appendix References. This change results in a new set of coefficient matrices. To calculate the attacked system’s square of voltages vector $(\textbf{U}^{A})$, we solve $\textbf{U}^{A}_{(N-1)\times 1}=\Omega^{A}_{(N-1)\times N}\begin{bmatrix}1\\ \textbf{U}^{A}\end{bmatrix}_{N\times 1}$. To obtain the coefficient matrices, (13) is dragged into LinDistFlow as the ZP model is imposed on them. The final results are $\omega_{i,k}^{A}=\omega_{i,k}+\omega_{i,k}^{a}$, for $i\geq 2$ and $k\geq 1$; $\Omega^{A}_{(N-1)\times N}=[\Omega^{A^{\prime\prime}}_{(N-1)\times 1}\ \Omega^{A^{\prime}}_{(N-1)\times(N-1)}]$, in which for $i\geq 2$ and $k\geq 2$ and $\omega_{i,1}^{a}=\sum_{c\in\{\mathcal{D}_{i}\cap\mathcal{D}_{a}\}}-2p_{a}^{A_{0}}\alpha^{\prime}_{p}r_{\pi_{c},c}-2q_{a}^{A_{0}}\alpha^{\prime}_{q}x_{\pi_{c},c}$ , $\omega_{i,k}^{a}=\begin{cases}-2r_{\pi_{i},i}p^{a}\gamma^{\prime}_{p}-2x_{\pi_{i},i}q^{a}\gamma^{\prime}_{q},\ \ \text{if}\hskip 5.69046pti\in\mathcal{D}_{a}\\ \omega_{\pi_{i},k}^{A},\ \ \text{otherwise}.\end{cases}$
@@
Note that when there is an attack in a leaf bus (the last bus of each branch), $r_{k,a}$ and $x_{k,a}$ have the highest possible values. As a result, the voltage drop resulted from $(-2p^{a}r_{k,a}-2q^{a}x_{k,a})$ in (17) is higher and obtained voltages shrink. In conclusion, attacks on the leaf buses yield the most detrimental effects. However, this conclusion only holds when there are no DERs in the system. Since DERs can be modeled as negative loads, the direction of power flow on some lines may change, and the voltage drop trend may not be the same.
@@ III-B Analytical Insights into the Attack Impact
Using the results obtained in Subsection III-A, we further obtain the minimum number of attacked devices, which leads to voltage safety violations. We call such a threat the “critical attack”. For this, we consider the voltage of the leaf bus as a known variable ($v_{th}$). The new unknown variable is $p^{a}$, and based on the attacked device type, we can find $q^{a}$ via $q^{a}=\frac{q_{d}}{p_{d}}\ p^{a}$. So the new set of coefficients for obtaining voltages of buses except for the leaf one and the active power of the critical attack is forming $\Omega^{d}$ ($\Omega^{d}=[\Omega^{d^{\prime\prime}}\Omega^{d^{\prime}}]$) in which
$\omega_{i,1}^{d}$ $=\begin{cases}\omega_{i,1}+v_{th}^{2}\omega_{i,a},\ \ \text{if}\ i\neq a,\\ \omega_{i,1}+v_{th}^{2}(\omega_{i,a}-1),\ \ \text{if}\ i=a,\end{cases}$ (18)
$\omega_{i,k}^{d}$ $=\begin{cases}\sum_{c\in\mathcal{D}_{i}}-2r_{\pi_{c},c}\alpha^{\prime}_{p}-2\frac{q_{d}}{p_{d}}x_{\pi_{c},c}\alpha^{\prime}_{q},\ \ \\ \quad\quad\quad\ \text{if}\ k=a,i\in\mathcal{D}_{a},\\ \omega_{\pi_{i},k}^{d},\ \ \text{if}\ k=a,i\notin\mathcal{D}_{a},\\ \omega_{i,k},\ \ \text{otherwise}.\end{cases}$ (19)
Hence, we can solve the linear system of equations given by $(\textbf{I}^{a}-\Omega^{d^{\prime}})\textbf{X}=\Omega^{d^{\prime\prime}}$, in which X is a vector with the same dimension as U where all elements of it are the same as U except for one in which X contains $p^{a}$ instead of $u_{a}$ (since we already know $u_{a}=v_{th}^{2}$). Additionally, $\textbf{I}^{a}$ is the identity matrix except for the element $(a,a)$ which equals 0. This gives the bus voltages as well as the required $p^{a}$. Then, we can find the number of required devices in bus $n$ using $c(n)=\frac{p^{a}}{p_{d}}$.
The closed-form expressions provided in this section are not only used to identify the worst effect of the LAAs, but also inspire our mitigation method to find the optimal defensive action in the following sections. Additionally, the introduced “critical attack” has been implemented in finding the optimal action of the strategic attacker later in this paper.